In a race of two cars, the first car travels at a speed of 20 km/hr whereas the second car travels at a speed of 18 km/hr. If the second car arrives 6 minutes later than the first car then, what is the distance of the start to destination?
Answer by richwmiller(17219)

20*t=18*(t+1/10)
20t=18t+1.8
2t=1.8
t=.9
check
20*.9=18*1
18=18
ok